# HCOA*: Hierarchical Class-ordered A* for Navigation in Semantic Environments

### v1

@@ VII Conclusions
In this paper, we addressed the problem of robot navigation in 3D geometric and semantic environments by leveraging the hierarchical structure of the environment for efficient path-planning. We introduced Hierarchical Class-ordered A* (HCOA*), an algorithm that exploits a total order over semantic classes to guide the search process while significantly reducing computational effort. To classify higher-layer nodes, we proposed two methods: a Graph Neural Network-based method and a Majority-Class method. Through simulations on a 3D Scene Graph, we showed that HCOA* effectively reduces computational cost compared to a state-of-the-art method. Specifically, our results show that HCOA* finds the optimal path while reducing the number of expanded nodes by 25% and achieving a 16% reduction in computational time for a typical 3DSG with 1,314 nodes.
Moving forward, future work will focus on applying HCOA* to larger, real-world datasets to demonstrate the full potential of the algorithm. Additionally, we aim to develop an online path-planning framework based on HCOA*, enabling adaptive decision-making in real-time, dynamic scenarios.

### v2

@@ VII Conclusion
In this paper, we have addressed the problem of robot navigation in 3D geometric/semantic environments by leveraging the environment’s hierarchy for efficient path-planning. We have introduced Hierarchical Class-ordered A* (HCOA*), an algorithm that exploits a total order over semantic classes to guide the search process while significantly reducing computational effort. To classify higher-layer nodes, we proposed three methods: a Majority-Class method, a k-Nearest Neighbors method, and a Graph Neural Network method. Through simulations on two 3D Scene Graphs, we showed that HCOA* effectively reduces computational cost compared to other state-of-the-art methods. Specifically, our results show that HCOA* reduces the computational time of navigation by up to 50%, while maintaining near-optimal performance.
Future work will address the limitations of using a total order over semantic classes for navigation. While total ordering offers a simple and interpretable mechanism for prioritizing semantic categories, it is insufficient for capturing more complex or temporally dependent constraints (e.g., avoid kitchen after visiting bathroom). To handle such cases, we plan to incorporate temporal logic [22] and explore partial ordering schemes. These enhancements will also necessitate extending our current path-planning framework to support online replanning capabilities. To this end, we aim to develop an online path-planning framework based on HCOA*, enabling adaptive decision-making in real-time, dynamic scenarios.